# Stable coresets: Unleashing the power of uniform sampling

### v1

@@ Appendix C Proofs omitted from Section 4
Let $\mu\in\mathbb{R}^{d}$ be a median of $P\subset\mathbb{R}^{d}$ of size $n=|P|$, and assume for simplicity that $n$ is odd. Without loss of generality we may assume that $P$ contains no repeated points, as duplicate points can be perturbed by an infinitesimal amount. We will also need the following well-known fact.
@@
We first consider $d=1$, and without loss of generality assume $x\leq\mu$. To evaluate the difference in average costs, we break the calculation into the three regions $L$, $M$, and $U$: $L:=\{p\in P:\ p\leq x\}$, $M:=\{p\in P:\ x<p\leq\mu\}$, and $U:=\{p\in P:\ \mu<p\}$. In each region, we can simplify the expression $|x-p|-|\mu-p|$ by expressing the absolute values explicitly, which allows us to evaluate $\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(\mu,P)$. Notably, in regions $L$ and $U$, the value of $|x-p|-|\mu-p|$ depends only on $x$ and $\mu$, not on the specific value of $p$. Observe that $\left\lvert L\right\rvert=\left\lvert P\right\rvert\operatorname{edf}_{P}(x)$, $\left\lvert M\right\rvert=\left\lvert P\right\rvert(\operatorname{edf}_{P}(\mu)-\operatorname{edf}_{P}(x))$ and $\left\lvert U\right\rvert=\left\lvert P\right\rvert(1-\operatorname{edf}_{P}(\mu))=\frac{|P|}{2}$. We can write
@@
[…]
@@
We now bound the term $\frac{1}{|Q|}\sum_{q\in M\cap Q}|\mu-q|-\frac{1}{|P|}\sum_{p\in M}|\mu-p|$. Partition $M$ to intervals $I_{0},I_{1},...,I_{t}$, such that each interval contains exactly $2\epsilon|P|$ points, except for the last one which might be smaller.
Let $a_{i}=\frac{1}{|Q|}|I_{i}\cap Q|-\frac{1}{|P|}|I_{i}\cap P|$ for i=0,,ti=0,...,t.
@@
Each interval contains at least 0 points from $Q$ and at most $4\epsilon|Q|$ points from $Q$. Consequently, $2\epsilon\leq a_{i}\leq 2\epsilon$. We can write
@@
where the last inequality follows from the fact that for every $j_{1}\leq j_{2}$ we have $2\epsilon\leq\sum_{i=j_{1}}^{j_{2}}a_{i}\leq 2\epsilon$.
@@ Appendix D Proofs omitted from Section 5
Let $P\subset\mathcal{S}_{d}$ be finite and let $0<\epsilon\leq\frac{1}{5}$. Then a uniform sample of size $O(\epsilon^{-2}\log d)$ from $P$ is a stable $(\epsilon/6,4\epsilon)$-coreset for $1$-median in $\mathcal{S}_{d}$ under the Kendall-tau metric with probability at least $4/5$.
@@ D.1 Low-distortion embeddings
$\forall p\in P,\forall y\in\mathcal{X}_{1},\qquad\tfrac{1}{D}\cdot\operatorname{dist}_{2}(f(x),f(y))\leq r\cdot\text{dist}_{1}(x,y)\leq D\cdot\operatorname{dist}_{2}(f(x),f(y)).$ (11)
@@ D.3 $C$-dispersed instances
As before, let $\mu$ denote a median of $P$. First note that if $P$ is a bounded instance then $\max_{x,y\in P}\left\lVert x-y\right\rVert_{1}\leq\frac{C}{|P|^{2}}\sum_{x,y\in P}\left\lVert x-y\right\rVert_{1}\leq\frac{2C}{|P|}\sum_{x\in P}\left\lVert\mu-x\right\rVert_{1}=2C\operatorname{\overline{opt}}(P)$. By the triangle inequality, for every $x\in P$,
@@
Additionally, we have that $\operatorname*{\mathbb{E}}[\operatorname{\overline{cost}}(\mu,Q)]=\operatorname{\overline{cost}}(\mu,P)=\operatorname{\overline{opt}}(P)$. Hence, using Hoeffding’s bound.
$\Pr\Big[\Big|\operatorname{\overline{cost}}(\mu,Q)-\operatorname{\overline{cost}}(\mu,P)\Big|<\frac{\epsilon}{2}\operatorname{\overline{cost}}(\mu,P)\Big]>1-2\exp{\frac{-\epsilon^{2}|Q|}{2(2C+1)^{2}}}>1-\frac{1}{d^{2}}.$
Combining the union bound and Theorem 1.4, with probability at least $\frac{3}{4}$, $Q$ is $\frac{\epsilon}{2}$-RCDA and for every $x\in\mathbb{R}^{d}$:
$\left\lvert\operatorname{\overline{cost}}(m,P)-\operatorname{\overline{cost}}(m,Q)\right\rvert\leq\frac{\epsilon}{2}\operatorname{\overline{cost}}(m,P).$
Consequently, we obtain that $Q$ is a strong $\epsilon$-coreset.
$\left\lvert\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(x,Q)\right\rvert$=
$\left\lvert\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(m,P)+\operatorname{\overline{cost}}(m,P)-\operatorname{\overline{cost}}(x,Q)+\operatorname{\overline{cost}}(m,Q)-\operatorname{\overline{cost}}(m,Q)\right\rvert\leq$
$\left\lvert[\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(m,P)]-[\operatorname{\overline{cost}}(x,Q)-\operatorname{\overline{cost}}(m,Q)]\right\rvert+\left\lvert\operatorname{\overline{cost}}(m,P)-\operatorname{\overline{cost}}(m,Q)\right\rvert$
$\leq\frac{\epsilon}{2}\operatorname{\overline{cost}}(x,P)+\frac{\epsilon}{2}\operatorname{\overline{cost}}(m,P)\leq\epsilon\operatorname{\overline{cost}}(x,P).$

### v2

@@ Appendix C Proofs omitted from Section 4
Let $P\subset\mathbb{R}^{d}$ of size $n=|P|$, and assume for simplicity that $n$ is odd. Without loss of generality, we may assume that $P$ contains no repeated points, as duplicate points can be perturbed by an infinitesimal amount. Denote by $\mu\in\mathbb{R}^{d}$ a median of $P$. We will also need the following well-known fact.
@@
We first consider $d=1$, and without loss of generality assume $x\leq\mu$. To evaluate the difference in average costs, we break the calculation into the three regions $L$, $M$, and $U$ (similar strategy was used in [Dan21]): $L:=\{p\in P:\ p\leq x\}$, $M:=\{p\in P:\ x<p\leq\mu\}$, and $U:=\{p\in P:\ \mu<p\}$. In each region, we can simplify the expression $|x-p|-|\mu-p|$ by expressing the absolute values explicitly, which allows us to evaluate $\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(\mu,P)$. Notably, in regions $L$ and $U$, the value of $|x-p|-|\mu-p|$ depends only on $x$ and $\mu$, not on the specific value of $p$. Observe that $\left\lvert L\right\rvert=\left\lvert P\right\rvert\operatorname{edf}_{P}(x)$, $\left\lvert M\right\rvert=\left\lvert P\right\rvert(\operatorname{edf}_{P}(\mu)-\operatorname{edf}_{P}(x))$ and $\left\lvert U\right\rvert=\left\lvert P\right\rvert(1-\operatorname{edf}_{P}(\mu))=\frac{|P|}{2}$. We can write
@@
Similarly for $Q$ (the fact that $\mu$ is a median of $P$ is not utilized in the argument above),
@@
We now bound the term $\frac{1}{|Q|}\sum_{q\in M\cap Q}|\mu-q|-\frac{1}{|P|}\sum_{p\in M}|\mu-p|$. Partition $M$ to intervals $I_{0},I_{1},...,I_{t}$, such that each interval contains exactly $2\epsilon|P|$ points, except for the last one which might be smaller. We denote $I_{i}=(p_{i},p_{i+1}]$, with $p_{0}=x$ and $p_{t+1}=\mu$ and $a_{i}=\frac{1}{|Q|}|I_{i}\cap Q|-\frac{1}{|P|}|I_{i}\cap P|$ for i=0,,ti=0,...,t.
@@
Each interval contains at least 0 points from $Q$ and at most $4\epsilon|Q|$ points from $Q$. Consequently, $2\epsilon\leq\sum_{i=j_{1}}^{j_{2}}a_{i}\leq 2\epsilon$, for every $j_{1}\leq j_{2}$. We can write
@@ Appendix D Proofs omitted from Section 5
Let $P\subset\mathcal{S}_{d}$ be finite and let $\epsilon\in(0,\frac{1}{5})$. Then a uniform sample of size $O(\epsilon^{-2}\log d)$ from $P$ is a stable $(\epsilon/6,4\epsilon)$-coreset for $1$-median in $\mathcal{S}_{d}$ under the Kendall-tau metric with probability at least $4/5$.
@@ D.1 Low-distortion embeddings
$\forall p\in P,\forall x\in\mathcal{X}_{1},\qquad\tfrac{1}{D}\cdot\operatorname{dist}_{2}(f(p),f(x))\leq r\cdot\text{dist}_{1}(p,x)\leq D\cdot\operatorname{dist}_{2}(f(p),f(x)).$ (11)
@@ D.3 $C$-dispersed instances
As before, let $\mu$ denote a median of $P$. By the triangle inequality,
$\max_{x,y\in P}\left\lVert x-y\right\rVert_{1}\leq\frac{C}{|P|^{2}}\sum_{x,y\in P}\left\lVert x-\mu+\mu-y\right\rVert_{1}\leq\frac{2C}{|P|}\sum_{x\in P}\left\lVert\mu-x\right\rVert_{1}=2C\operatorname{\overline{opt}}(P),$
and thus for every $x\in P$,
@@
Additionally, $\operatorname*{\mathbb{E}}[\operatorname{\overline{cost}}(\mu,Q)]=\operatorname{\overline{cost}}(\mu,P)=\operatorname{\overline{opt}}(P)$, and thus Chernoff’s bound implies
$\Pr\Big[\Big|\operatorname{\overline{cost}}(\mu,Q)-\operatorname{\overline{cost}}(\mu,P)\Big|\geq\frac{\epsilon}{2}\operatorname{\overline{cost}}(\mu,P)\Big]\leq 2\exp\left(-\frac{\epsilon^{2}|Q|}{3(2C+1)}\right)\leq\frac{\delta}{2}.$
Following Corollary 4.5 and Lemma 4.6, with probability at least $1-\frac{\delta}{2}$, the subset $Q$ is $\frac{\epsilon}{2}$-RCDA. By the union bound, with probability at least $1-\delta$, we have that $Q$ is $\frac{\epsilon}{2}$-RCDA and $\left\lvert\operatorname{\overline{cost}}(\mu,P)-\operatorname{\overline{cost}}(\mu,Q)\right\rvert\leq\frac{\epsilon}{2}\operatorname{\overline{cost}}(\mu,P)$. In this case, for every $x\in\mathbb{R}^{d}$,
$\left\lvert\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(x,Q)\right\rvert$
$\leq\left\lvert[\operatorname{\overline{cost}}(x,P)-\operatorname{\overline{cost}}(\mu,P)]-[\operatorname{\overline{cost}}(x,Q)-\operatorname{\overline{cost}}(\mu,Q)]\right\rvert+\left\lvert\operatorname{\overline{cost}}(\mu,P)-\operatorname{\overline{cost}}(\mu,Q)\right\rvert$
$\leq\frac{\epsilon}{2}\operatorname{\overline{cost}}(x,P)+\frac{\epsilon}{2}\operatorname{\overline{cost}}(\mu,P)\leq\epsilon\operatorname{\overline{cost}}(x,P),$
which establishes that $Q$ is a strong $\epsilon$-coreset. ∎